The average speed of an airplane is 10 times that of a car. It takes the airplane 18h less than the car to travel 1000km. Determine the average speeds of the air plane and the car.
Answer by mananth(16946)

car speed =x
plane speed = 10x
distance =1000 km
time taken by air plane = 1000/10x
time taken by car = 1000/x
Time taken by car - time taken by plane = 18
1000/x - 1000/10x= 18
LCD = 10x
multiply equation by 10x
10000 -1000 = 18(10x)
9000 = 180x
/180
x= 9000/180
x= 50
car speed = 50 km/h
plane speed = 10x -> 10*50 = 500 km/h